Present: Rich Bennett, Cynthia Bowker, Will Chaney, Bill Covey, Suzy Covey, Debra Harris, Winston Harris, Bobbie Meng, Hilton Soto, Geoffrey Stewart, Ying Tang
Ongoing Systems Updates
- Geoffrey is getting a new motorcycle – a BMW R100S.
- Kinetic art update: it is actually sending bad Morse code. Only one of the codes is correct. Suzy had five guys show up in her office today – looking for a place to put more pigeons.
- Bobbie’s stepdaughter placed 7th in weigh lifting competition this weekend. Her school came in 4th overall.
- Network File Sharing Basics – reviewed the procedures for setting up and accessing shared permission files.
- Kangaroo USB drives – are for Access Services’ document delivery program. They are 128M devices, solid state - that will enable file moving. They plan on scanning items in LAD, putting them on the drives, then carrying the drives back here to upload on the Web. The drives are very small – with no moving parts - need to come up with a way to identify & keep track of them.
- The next order for computers should go out soon.
- A back-up server for NT2R should be ordered next. A back-up server site has been located in room 206 MSL. The machine will be smaller and not clustered. No tape drives. We’ll want to get a locking cabinet.
- Scanner problem in DLC – Microteks – driver software was not coping with W2K. We replaced the SCSI cards with 2906’s – the 1505’s are no longer made. Scanners always need to powered-up first. The Polaroid transparency scanner is not working with it. It may need a new driver.
Software Projects Status
- Setting up an ILL/Access Services content site – we’re souping up a workstation for this – will keep it isolated and see how it works.
- Individual mailbox backup – last time Will checked, it had run fine. He is still removing corruption and now working on permission problems. Marcia Pearce’s folder is still a problem.
- Jeffrey Barr thought he had backed up email from 12/24/01-1/14/02, before he received his new machine - but apparently he had not. His old machine has been imaged, there’s nothing to restore. Will check on NT2.
- Anchor – Prototype is done. Will demo to Susan and John this week.
- Two new DLC projects – University Archives & Ariel. Both are largely photographs in a Web-based database. Ying will be working on both.
- Film Log – Problems with the old Film Log – it was getting an out of disk space message on NT4, but there was plenty of space there. Winston emptied the garbage on the C: which fixed it. Bill suggested checking the paging file, cut it, then put another paging file onto a different drive.
- ILL Document Delivery – needs an application to generate file names and passwords.
- Personnel – wants an extra page on their reports. Web Projects
- Migration to new server in process – may be ready to turn the new version over to a small group for testing at first – then over to a larger group if that is successful. Virtual directories are being created on some software programs.
Back to the Systems Homepage
Last updated February 4, 2002
by Debra Harris